# for Human Motion Generation

### v2

@@ 3.3 Motion Generator Fine-tuning
Existing motion generators are often trained with limited real motion data, which may limit their capabilities. However, collecting large-scale real 3D motion data is costly. To overcome this challenge, we leverage a large volume of synthetic, physically-plausible motion data provided by MPR module, to further enhance the capabilities of the motion generator.
@@ 4 Experiment
Extensive experiments evaluate the performance of our Morph across multiple motion generation tasks and datasets. Specifically, we assess Morph on two motion generation tasks, text-to-motion and music-to-dance. As Morph is agnostic to specific instantiation of generation models, we combine Morph with three types of generation models for text-to-motion, i.e., diffusion-based models, including MotionDiffuse [40] and MDM [31], autoregressive-based models, including T2M-GPT [40] and generative make modeling, including MoMask [7]. For music-to-dance, we combine Morph with the diffusion-based model EDGE [33] and autoregressive model Bailando [29]. The Appendix provides additional visualizations and experimental results showcasing Morph’s performance on varying amounts of noisy motion data, different threshold $\tau$ in imitation selection opeartion, and multi-round optimization of the MPR module and motion generator.
Table 1: Ablation study on Morph-MoMask (combined with MoMask [7] generator) for text-to-motion task on HumanML3D dataset. IS: imitation selection operation; Adversarial: using adversarial reward training MPR module; Energy: using energy reward training MPR module; Real Data: using real motion data training MPR module; FT: fine-tuning Motion Generator with physics-refined motions. The arrows ($\uparrow$/$\downarrow$ / ↓) indicate that higher/smaller values are better.
@@
IS Adversarial Energy Real Data FT RTOP-3 $\uparrow$ FID $\downarrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
@@
$\mathrm{B}$ 0.792 0.194 0.852 0.000 2.272 0.020 0.0155
$\mathrm{C}$ 0.782 0.276 0.735 0.000 2.554 0.032 0.0362
$\mathrm{D}$ 0.790 0.165 0.715 0.000 2.376 0.026 0.0272
$\mathrm{E}$ 0.801 0.077 0.672 0.000 2.271 0.011 0.0155
$\mathrm{F}$ 0.785 0.183 0.749 0.000 2.451 0.017 0.0338
$\mathrm{G}$ 0.816 0.043 0.651 0.000 2.146 0.010 0.0151
@@ 4.1 Datasets and Evaluation Metrics
Evaluation Metrics. We evaluate motion generation from two perspectives: common generation metrics and physical plausibility metrics. We first present the common generation metrics. For text-to-motion generation, following [8, 41], we use three standard metrics: Frechet Inception Distance (FID) that measures the distance between the generated and ground-truth motion distributions; R-Precition (including RTOP-1, RTOP2, and RTOP-3) that computes the retrieval accuracy of the generated motions with respect to the input text; Diversity (Div) that measures the diversity of generated motions. For music-to-dance generation, following [16, 38, 29], we use five standard metrics: FIDk and Divk that calculate Frechet Inception Distance and Diversity on kinetic features (denoted as ‘$k$’); FIDg and Divg that calculate Frechet Inception Distance and Diversity on geometric features (denoted as ‘$g$’); Beat Align Score (BAS) that measures the alignment between input music and generated dances.
@@ 4.2 Implementation Details
Training Setup. We implement Morph based on PyTorch. The NVIDIA’s Isaac Gym is used as the physics simulator. Rather than using real motion data, we train Morph on synthetic motion data produced by a pre-trained model, increasing the motion data volume to three times the original. In the training phase of MPR module (Stage 1), we use the Adam optimizer with a batch size of $64$ and a learning rate of $4\times 10^{-5}$. During the finetuning of motion generator (Stage 2), we follow the original training setup, only adjusting the learning rate to $1\times 10^{-5}$. For hyper-parameters setting, we set $w_{\theta}$, $w_{q}$, $w_{v}$ and $w_{w}$ (Eq. 2) to $0.5$, $0.3$, $0.1$ and $0.1$, $\alpha_{\theta}$, $\alpha_{q}$, $\alpha_{v}$ and $\alpha_{w}$ (Eq. 2) to $100$, $10$, $0.1$ and $0.1$, $\gamma$ (Eq. 5) is set to $0.1$ and the imitation selection threshold $\tau$ is set to $0.5$. All our experiments are conducted on 8 Tesla V100 GPUs. Further details on implementation and hyperparameter analysis are provided in the Appendix.
Data Preprocessing. In text-to-motion generation, some motion sequences involve environmental interactions, such as sitting down, climbing stairs, and swimming, which cannot be simulated by current physical simulator. Therefore, we annotate each text description in HumanML3D [9] based on its semantic content. Text Descriptions involving interaction-based motions are labeled as ‘00’, while other descriptions are labeled as ‘$1$’. Text labeled as ‘$1$’ will be used to train the MPR module in Stage 1. We will release this annotation on HumanML3D in the final version.
@@
Methods RTOP-1 $\uparrow$ RTOP-2 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ MM-Dist $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
@@
Morph-MDM 0.467 0.658 0.752 0.485 3.106 9.863 2.177
Morph-MD 0.494 0.687 0.786 0.571 3.001 9.492 1.584
Morph-T2M-GPT 0.497 0.684 0.781 0.109 2.997 9.768 1.903
Morph-MoMask 0.525 0.722 0.816 0.043 2.923 9.687 1.652
@@
Morph-Bailando 26.32 9.43 7.86 6.55 0.2411 0.045 0.000 2.061 0.023 0.0059
The generated motion sequences may exhibit issues such as body leaning, floating and ground penetration. When imported into the engine, these issues can cause instability in the robot, leading to falls, bouncing off the ground, or dropping from mid-air. To address this, we apply a preprocessing step to the motion sequences. Specifically, we first compute the body’s tilt angle, defined as the angle between the projection of the center of mass onto the ground and the line connecting both feet, and apply this angle to the pelvis for the entire sequence. To correct for floating and penetration, we determine the lowest point height and adjust the entire sequence by this offset. The resulting sequence is then used for training and inference. More details can be found in the Appendix.
@@ 4.3 Ablation Studies
Effectiveness of training MPR module using only generated data. We investigate the effectiveness of using only synthetic motion data from two perspectives: its overall effectiveness and its performance relative to training with real motion data. First, we compare the baseline model $\mathrm{A}$ (only the MoMask generator) and model $\mathrm{E}$ which combines MoMask with the trained MPR module. As shown in Tab. 1, model $\mathrm{E}$ outperforms model $\mathrm{A}$ on all physical metrics ($0.651/1.058$ on PFC, $0.0/23.152$ on Penetrate, $2.146/10.66$ on Float, $0.01/5.262$ on Skate), while maintaining competitive performance on generation metrics. These results confirm the feasibility of training the MPR module solely with synthetic data. Next, we compare model $\mathrm{E}$ with model $\mathrm{F}$, which uses real motion data to train the MPR module. As shown in Tab. 1, model $\mathrm{F}$ achieves inferior performance compared to model $\mathrm{E}$. This can be attributed to the significant domain gap between the training data (real motions) and test data (generated motions). This gap hinders the MPR module trained on real data from effectively adapting to the generated motion data, leading to performance decline.
@@
Effectiveness of adversarial reward and energy reward. When training the MPR module, we add an adversarial reward and energy reward in addition to the common imitation task reward. In Tab. 1, comparing model $\mathrm{E}$ with model $\mathrm{C}$, which omits the adversarial reward, we observe a performance decline in both generation and physical metrics on model $\mathrm{C}$. The significant drop in FID metric indicates that, without the adversarial reward, the simulated motions lack the distribution constraints of the input motions, leading to a deviation from the intended motion distribution. Next, we compare model $\mathrm{E}$ with $\mathrm{D}$, which omits the energy reward. As shown in Tab. 1, model $\mathrm{E}$ outperforms model $\mathrm{D}$ in generation metrics. We argue that introducing energy reward helps suppress high-frequency jitter during the humanoid robot control process, resulting in more natural motions.
@@
Effectiveness of imitation selection strategy. Like [39, 11], the proposed Morph focuses on addressing physical inconsistencies in non-interactive and ground-contact human motions. In training and testing data, there are some interactive and non-grounded motions, such as sitting on a chair, swimming, or climbing stairs. To ensure the quality of the motions after refining, we design an imitation selection operation to filter out the simulations of such motions. In Tab. 1, we compare model $\mathrm{E}$ with $\mathrm{B}$, which omits the imitation selection operation. As shown, model $\mathrm{E}$ outperforms $\mathrm{B}$ in both generation and physic metrics, validating the effectiveness of the imitation selection strategy.
@@
Effectiveness of finetuning motion generator using simulated motions. Furthermore, we compare model $\mathrm{E}$ and model $\mathrm{G}$, which finetunes the motion generator using the physically refined motions produced by the MPR module. As shown in Tab. 1, model $\mathrm{G}$ achieves further gains in both generation and physic metrics, with RTOP-3 increasing by $1.5\%$ and FID improving by $0.034$. These results suggest that the proposed MPR module can, in turn, enhance the performance of the motion generator.
@@ 4.4 Evaluation with Different Motion Generators
We evaluate the adaptability of our Morph with different motion generators, including MDM [31], Motiondiffuse [41], T2M-GPT [40], and MoMask [7]. As shown in Tab. 2, Morph significantly enhances physical fidelity while maintaining competitive generation metrics across different motion generation models. For example, after integrating Morph with MDM (Morph-MDM), RTOP-1 metric improved by $0.012$. In terms of physical plausibility metrics, penetration dropped to zero, float decreased from $17.502$ to $2.261$, and skate reduced from $3.540$ to $0.018$. Similar performance improvements are observed with other motion generators. The consistent gains across different generators highlight the versatility of our Morph framework.
@@ 4.5 Comparisons with State-of-the-arts
Main results on Music-to-Dance Generation. On the music-to-dance dataset, AIST++, we compare Morph combined with EDGE [33] and Bailando [29] generators against other state-of-the-art methods. As shown in Tab. 4, Morph significantly improves physical metrics compared to existing methods. In terms of generation metrics, our Morph-EDGE model achieves the best results in multiple metrics, including FIDg, Divg, and BAS, while also demonstrating competitive performance in FIDk metric. These results validate the superiority of our framework in music-to-dance generation.
@@ 5 Conclusion
In this paper, we present Morph, a model-agnostic physical optimization framework, designed to enhance physical plausibility in motion generation without relying on costly real-world motion data. To accomplish this, we first leverage a pretrained motion generator to synthesize large-scale noisy motion data. We then introduce a Motion Physics Refinement module, which utilizes these synthesized data to train a motion imitator that enforces physical constraints. The physically refined motions can, in turn, be used to finetune the motion generator, further enhancing its capabilities. Our framework is compatible with various motion generation models across both text-to-motion and music-to-dance generation tasks. Extensive experimental results demonstrate that Morph substantially improves physical plausibility while achieving competitive generation quality.
@@
In this Appendix, we present more details for Morph, including data preprocess, additional experimental results, qualitative comparisons. First, we describe the data preprocessing procedure used for training the Motion Physics Refinement (MPR) module with generated motion data (Sec. A). Then, we present experimental results analyzing the impact of $\tau$ in the imitation selection operation (Sec. B), the effect of varying the quantity of noisy motion data for MPR training (Sec. C), and effect of the number of training rounds for Morph (Sec. D). Finally, we provide additional qualitative comparisons for text-to-motion and music-to-dance tasks (Sec. E).
@@ Appendix A Details for Data Preprocess
One-Round w/o FT 0.511 0.801 0.077 9.579 0.672 0.000 2.271 0.011 0.0155
One-Round 0.525 0.816 0.043 9.687 0.651 0.000 2.146 0.010 0.0151
Two-Round w/o FT 0.525 0.818 0.043 9.689 0.635 0.000 2.132 0.010 0.0137
Two-Round 0.527 0.821 0.044 9.692 0.631 0.000 2.117 0.008 0.0134
@@ Appendix D Effect of Multi-Round Optimization of the MPR module and MG on Morph
As shown in Tab. 7, in the first round of training, MG improves the performance of MPR module, enhancing the physical quality of its generated motion. The refined motion data from the trained MPR module is then used to fine-tune the MG, boosting its performance further. In the second round, the fine-tuned MG from the first round is used to generate training data for the MPR module (initialized with first-round weights). We observed improvements in Two-Round w/o FT compared to One-Round, with PFC increasing by 0.016, Float by 0.014, and IFR decreasing, indicating enhanced motion imitation by the MPR module. After fine-tuning the MG once again, Two-Round shows improvements in the RTOP-1 and RTOP-3 metrics. These results clearly demonstrate that the MG and MPR modules can mutually enhance each other. Moreover, alternating training between the MG and MPR modules across multiple rounds can further improve the performance of Morph.
Appendix E More Qualitative Results

### v3

@@ 3.3 Motion Generator Fine-tuning
Existing motion generators are often trained with limited real motion data, which may limit their capabilities. However, collecting large-scale real 3D motion data is costly. So, we leverage a large volume of synthetic, physically-plausible motion data provided by MPR module, to further enhance the capabilities of the motion generator.
@@
Multi-Round Iterative Optimization.
In Fig. 2, we define the complete two-stage process as a single training round. To maximize the synergy between the MG and MPR modules, we repeat this process multiple times, using the model weights from each round as initialization for the next, thus creating a multi-round iterative training paradigm. We present the results of this iterative training in the Appendix, demonstrating its effectiveness and potential.
@@ 4 Experiment
Extensive experiments evaluate the performance of our Morph across multiple motion generation tasks and datasets. Specifically, we assess Morph on two motion generation tasks, text-to-motion and music-to-dance. As Morph is agnostic to specific instantiation of generation models, we combine Morph with three types of generation models for text-to-motion, i.e., diffusion-based models, including MotionDiffuse [42] and MDM [33], autoregressive-based models, including T2M-GPT [42] and generative make modeling, including MoMask [6]. For music-to-dance, we combine Morph with the diffusion-based model EDGE [35] and autoregressive model Bailando [31]. The Appendix provides additional visualizations and experimental results showcasing Morph’s performance on varying amounts of noisy motion data, different and threshold $\tau$ in imitation selection opeartion, and multi-round optimization of the MPR module and motion generator.
Table 1: Ablation study on Morph-MoMask (combined with MoMask [6] generator) for text-to-motion task on HumanML3D dataset. IS: imitation selection operation; Prior: using prior reward training MPR module; Adv: using adversarial reward training MPR module; Energy: using energy reward training MPR module; Real Data: using real motion data training MPR module; FT: fine-tuning Motion Generator with physics-refined motions. The arrows ($\uparrow/\downarrow$) indicate that higher/smaller values are better.
@@
IS Prior Adv Energy Real Data FT RTOP-3 $\uparrow$ FID $\downarrow$ PFC $\downarrow$ Penetrate $\downarrow$ Float $\downarrow$ Skate $\downarrow$ IFR $\downarrow$
@@
$\mathrm{B}$ 0.800 0.172 0.821 0.000 2.233 0.016 0.0152
$\mathrm{C}$ 0.792 0.194 0.852 0.000 2.272 0.020 0.0155
$\mathrm{D}$ 0.782 0.276 0.735 0.000 2.554 0.032 0.0362
$\mathrm{E}$ 0.795 0.163 0.711 0.000 2.362 0.024 0.0270
$\mathrm{F}$ 0.802 0.074 0.669 0.000 2.268 0.011 0.0153
$\mathrm{G}$ 0.792 0.179 0.734 0.000 2.420 0.015 0.0327
$\mathrm{H}$ 0.816 0.041 0.647 0.000 2.141 0.010 0.0149
@@ 4.1 Datasets and Evaluation Metrics
Evaluation Metrics. We evaluate motion generation from two perspectives: common generation metrics and physical plausibility metrics. First, for text-to-motion generation, we follow [7, 43] and use three standard metrics: Frechet Inception Distance (FID), which measures the distance between generated and ground-truth motion distributions; R-Precision (RTOP-1, RTOP-2, RTOP-3), which computes retrieval accuracy of generated motions based on input text; and Diversity (Div), which measures the diversity of generated motions. For music-to-dance generation, following [15, 40, 31], we use five metrics: FIDk and Divk for kinetic features (denoted ‘$k$’); FIDg and Divg for geometric features (denoted ‘$g$’); and Beat Align Score (BAS), which measures the alignment between input music and generated dances.
@@ 4.2 Implementation Details
Training Setup. We implement Morph based on PyTorch. The NVIDIA’s Isaac Gym is used as the physics simulator. During the MPR module’s training phase (Stage 1), we employ the Adam optimizer, with a batch size of 64 and a learning rate of $4\times 10^{-5}$. During the finetuning of motion generator (Stage 2), we follow the original training setup, only adjusting the learning rate to $1\times 10^{-5}$. For hyper-parameters setting, we set $w_{\theta}$, $w_{q}$, $w_{v}$ and $w_{w}$ (Eq. 2) to $0.5$, $0.3$, $0.1$ and $0.1$, $\alpha_{\theta}$, $\alpha_{q}$, $\alpha_{v}$ and $\alpha_{w}$ (Eq. 2) to $100$, $10$, $0.1$ and $0.1$ and the imitation selection threshold $\tau$ is set to $0.5$. All our experiments are conducted on 8 Tesla V100 GPUs. Further details on implementation and hyperparameter analysis are provided in the Appendix.
Data Preprocessing. In text-to-motion generation, some motion sequences involve environmental interactions, such as sitting down, climbing stairs, and swimming, which cannot be simulated by current physical simulator. Therefore, we annotate each text description in HumanML3D [8] based on its semantic content. Text Descriptions involving interaction-based motions are labeled as ‘0’, while other descriptions are labeled as ‘$1$’. Text labeled as ‘$1$’ will be used to train the MPR module in Stage 1. We will release this annotation on HumanML3D in the final version.
@@
Methods RTOP-1 $\uparrow$ RTOP-2 $\uparrow$ RTOP-3 $\uparrow$ FID $\downarrow$ $M$-Dist $\downarrow$ Diversity $\uparrow$ MModality $\uparrow$
@@
Morph-MDM 0.471 0.664 0.757 0.482 3.103 9.865 2.179
Morph-MD 0.494 0.691 0.788 0.568 3.000 9.494 1.582
Morph-T2M-GPT 0.498 0.690 0.784 0.105 2.994 9.771 1.910
Morph-MoMask 0.525 0.726 0.816 0.041 2.920 9.689 1.655
@@
Morph-Bailando 26.25 9.38 7.88 6.62 0.2412 0.043 0.000 2.057 0.021 0.0055
@@ 4.3 Ablation Studies
Effectiveness of training MPR module using only generated data. We investigate the effectiveness of using only synthetic motion data from two perspectives: its overall effectiveness and its performance relative to training with real motion data. First, we compare the baseline model $\mathrm{A}$ (only the MoMask generator) and model $\mathrm{F}$ which combines MoMask with the trained MPR module. In Tab. 1, model $\mathrm{F}$ outperforms model $\mathrm{A}$ on all physical metrics ($0.669/1.058$ on PFC, $0.0/23.152$ on Penetrate, $2.268/10.66$ on Float, $0.011/5.262$ on Skate), while maintaining competitive performance on generation metrics. These results confirm the feasibility of training the MPR module solely with synthetic data. Next, we compare model $\mathrm{F}$ with model $\mathrm{G}$, which uses real motion data to train the MPR module. In Tab. 1, model $\mathrm{G}$ achieves inferior performance compared to model $\mathrm{F}$. This can be attributed to the significant domain gap between the training data (real motions) and test data (generated motions). This gap hinders the MPR module trained on real data from effectively adapting to the generated motion data, leading to performance decline.
@@
Effectiveness of prior reward and energy reward. When training the MPR module, we add an prior reward and energy reward alongside the common imitation task reward. In Tab. 1, comparing model $\mathrm{F}$ with model $\mathrm{D}$, which omits the prior reward, we observe a performance decline in both generation and physical metrics on model $\mathrm{D}$. The significant drop in FID metric indicates that, without the prior reward, the simulated motions lack the distribution constraints of the input motions, leading to a deviation from the intended motion distribution. The $\mathrm{B}$, which leverages prior rewards, significantly outperforms model $\mathrm{C}$, which uses adversarial rewards, across various generation and physics-based metrics. This highlights the superior effectiveness of the proposed prior reward over the adversarial reward. We compare model $\mathrm{F}$ with $\mathrm{E}$, which omits the energy reward. As shown in Tab. 1, model $\mathrm{F}$ outperforms model $\mathrm{E}$ in generation metrics. We argue that introducing energy reward helps suppress high-frequency jitter during the humanoid robot control process, resulting in more natural motions.
@@
Effectiveness of imitation selection strategy. Like [41, 10], the proposed Morph focuses on addressing physical inconsistencies in non-interactive and ground-contact human motions. In training and testing data, there are some interactive and non-grounded motions, such as sitting on a chair, swimming, or climbing stairs. To ensure the quality of the motions after refining, we design an imitation selection operation to filter out the simulations of such motions. In Tab. 1, we compare model $\mathrm{F}$ with $\mathrm{B}$, which omits the imitation selection operation. As shown, model $\mathrm{F}$ outperforms $\mathrm{B}$ in both generation and physic metrics, validating the effectiveness of the imitation selection strategy.
@@
Effectiveness of finetuning motion generator using simulated motions. Furthermore, we compare model $\mathrm{F}$ and model $\mathrm{H}$, which finetunes the motion generator using the physically refined motions produced by the MPR module. As shown in Tab. 1, model $\mathrm{H}$ achieves further gains in both generation and physic metrics, with RTOP-3 increasing by $1.4\%$ and FID improving by $0.033$. These results suggest that the proposed MPR module can, in turn, enhance the performance of the motion generator.
@@ 4.4 Evaluation with Different Motion Generators
We evaluate the adaptability of our Morph with different motion generators, including MDM [33], Motiondiffuse [43], T2M-GPT [42], and MoMask [6]. As shown in Tab. 2, Morph significantly enhances physical fidelity while maintaining competitive generation metrics across different motion generation models. For example, after integrating Morph with MDM (Morph-MDM), RTOP-1 metric improved by $0.016$. In terms of physical plausibility metrics, penetration dropped to zero, float decreased from $17.502$ to $2.258$, and skate reduced from $3.540$ to $0.016$. Similar performance improvements are observed with other motion generators. The consistent gains across different generators highlight the versatility of our Morph framework.
@@ 4.5 Comparisons with State-of-the-arts
Main results on Music-to-Dance Generation. On the music-to-dance dataset, AIST++, we compare Morph combined with EDGE [35] and Bailando [31] generators against other state-of-the-art methods. As shown in Tab. 4, Morph significantly improves physical metrics compared to existing methods. For generation metrics, our Morph-EDGE model achieves the best results in multiple metrics, including FIDg, Divg, and BAS, while also demonstrating competitive performance in FIDk. These results validate the superiority of our framework in music-to-dance generation.
@@ 5 Conclusion
In this paper, we present Morph, a model-agnostic physical optimization framework designed to enhance physical plausibility in motion generation without relying on costly real-world data. We begin by using a pretrained motion generator (MG) to synthesize large-scale noisy motion data. Then, we introduce a Motion Physics Refinement (MPR) module that leverages this synthesized data to train a motion imitator enforcing physical constraints. These physically refined motions are used to fine-tune the motion generator, further enhancing its capabilities. To stabilize the physics optimization process, we introduce a prior reward module. This collaborative training paradigm enables mutual enhancement between MG and MPR module, improving practicality and robustness. Our framework is compatible with various motion generation models for both text-to-motion and music-to-dance tasks. Extensive experiments show that Morph substantially improves physical plausibility while achieving competitive generation quality.
@@
In this Appendix, we present more details for Morph, including data preprocess, additional experimental results, qualitative comparisons. First, we describe the data preprocessing procedure used for training the Motion Physics Refinement (MPR) module with generated motion data (Sec. A). Then, we present experimental results analyzing the impact of $\tau$ in the imitation selection operation (Sec. B), the effect of varying the quantity of noisy motion data for MPR training (Sec. C), and effect of the number of training rounds for Morph (Sec. E). Finally, we provide additional qualitative comparisons for text-to-motion and music-to-dance tasks (Sec. H).
@@ Appendix A Details for Data Preprocess
One-Round w/o FT 0.516 0.802 0.074 9.578 0.669 0.000 2.268 0.011 0.0153
One-Round 0.525 0.816 0.041 9.689 0.645 0.000 2.141 0.010 0.0149
Two-Round w/o FT 0.526 0.817 0.041 9.692 0.632 0.000 2.129 0.010 0.0134
Two-Round 0.527 0.818 0.038 9.697 0.625 0.000 2.108 0.007 0.0129
Three-Round w/o FT 0.526 0.821 0.040 9.701 0.621 0.000 2.121 0.009 0.0131
Three-Round 0.528 0.823 0.034 9.715 0.618 0.000 2.100 0.006 0.0122
@@ Appendix C Effect of Varying Amounts of Noisy Motion Data on Morph
Figure 5: T-sne of motion and text distribution between MG and Morph.
Table 8: The win rate of Morph over baselines.
Module Semantic Consistency Realism Physical Plausibility Fluency
MDM-Morph vs. MDM 84.8% 80.1% 96.6% 85.0%
T2M-GPT-Morph vs. T2M-GPT 87.1% 79.6% 94.4% 81.2%
MoMask-Morph vs. MoMask 90.4% 88.3% 97.5% 80.9%
@@ Appendix D Effect of Multi-Round Optimization of the MPR module and MG on Morph
  • Three-Round w/o FT: Training the MPR module again using the motion data generated by the fine-tuned MG from the second round.
  • Three-Round: Fine-tuning the Motion Generator using the results from Three-Round w/o FT.
As shown in Tab. 7, in the first round of training, MG improves the performance of MPR module, enhancing the physical quality of its generated motion. The refined motion data from the trained MPR module is then used to fine-tune the MG, boosting its performance further. In the second round, the fine-tuned MG from the first round is used to generate training data for the MPR module (initialized with first-round weights). We observed improvements in Two-Round w/o FT compared to One-Round, with PFC increasing by 0.013, Float by 0.012, and IFR decreasing, indicating enhanced motion imitation by the MPR module. After fine-tuning the MG once again, Two-Round shows improvements in the RTOP-1 and RTOP-3 metrics. The model’s generation and physical performance reached their best in Three-Round. These results clearly demonstrate that the MG and MPR modules can mutually enhance each other. Moreover, alternating training between the MG and MPR modules across multiple rounds can further improve the performance of Morph.
Table 9: Cross-Task generalization results on Music2Dance and Text2Motion
Task $FID/FID_{k}$ $\downarrow$ $FID_{g}$ $\downarrow$ $RTOP3/Div_{k}$ $\uparrow$ PFCPFC $\downarrow$ PenetratePenetrate $\downarrow$ FloatFloat $\downarrow$
Cross-task generalization evaluation
Text-to-Motion (MPR trained on T2M) 0.074 0.802 0.669 0.000 2.268
Text-to-Motion (MPR trained on M2D) 0.116 0.795 0.881 0.000 2.436
Music-to-Dance (MPR trained on M2D) 35.48 7.70 12.03 0.044 0.000 2.076
Music-to-Dance (MPR trained on T2M) 37.66 7.38 14.02 0.057 0.000 2.193
Action-to-motion tests
Action_Label-to-Motion (MDM-action) 0.497 0.396 0.544 15.770 7.467
Action_Label-to-Motion (MDM-action with Morph) 0.424 0.416 0.509 0.000 2.115
GAN-based Transformer tests
Text-to-Motion (GAN-based Transformer) 0.628 - 0.736 0.933 47.612 21.008
Text-to-Motion (GAN-based Transformer with Morph) 0.606 - 0.755 0.742 0.000 2.637
Long-duration dance samples tests
Music-to-Dance (30s long-term dance, Lodge) 45.56 34.29 6.75 0.114 46.772 29.857
Music-to-Dance (30s long-term dance, Lodge-Morph) 43.96 32.88 6.90 0.083 0.000 3.163
PHC-based baseline
Text-to-Motion (MoMask+PHC) 0.183 - 0.785 0.749 0.000 2.451
Text-to-Motion (MoMask+Morph) 0.041 - 0.816 0.647 0.000 2.141
Appendix E Semantic Alignment Analysis
As shown in Fig. 5, Morph significantly outperforms MG on alignment metrics. Globally, Morph demonstrates superior realism by closely matching the statistical distribution of real motions—exhibiting similar clustering patterns and range of variation (Left). Crucially, it also achieves stronger semantic alignment with input text features, forming tighter clusters around corresponding text embeddings to better capture intended meanings (Middle). Locally, Morph provides enhanced semantic matching at the segment level, ensuring fine-grained motion elements correspond more accurately to detailed text semantics throughout the sequence (Right). In summary, Morph can generate semantically faithful, realistic motions compared to the MG baseline.
Appendix F Cross-Task Generalization Ability
To evaluate the cross-task generalization of the MPR module, we conducted cross-validation by testing its performance across two distinct tasks: text-to-motion (using the MoMask dataset) and music-to-dance (using Bailando). These tasks differ significantly in input modalities—one driven by linguistic descriptions, the other by rhythmic audio-and in motion characteristics, from daily actions to stylized dance moves. As shown in Tab. 9, the MPR module retains strong performance even when trained without task-specific synthetic data: it not only preserves motion quality (e.g., smooth transitions and natural postures) but also maintains physical plausibility (avoiding joint distortions or gravity-defying movements). This confirms its ability to generalize beyond specific task boundaries.
Appendix G User Study
In the user study, we evaluated win rates across four critical dimensions—semantic alignment (matching text descriptions), authenticity (resembling real motions), physical plausibility (avoiding unnatural joint movements), and fluency (smooth temporal transitions). Morph decisively outperformed baseline methods here: as shown in Tab. 8, it achieved significantly higher win rates of 87.4%, 85.2%, 96.2%, and 82.4% respectively. Such consistent leads across all key metrics confirm its comprehensive advantages in motion generation quality.
Appendix H More Qualitative Results